OpenStudy (anonymous):

Reduce the following rational expression to lowest terms, if possible. (Problem in comments)

OpenStudy (anonymous):

\[\frac{ 4m^3-4m^2-80m }{ m^2-7m+10 }\]

hartnn (hartnn):

you can factor out 4m from the numerator

OpenStudy (anonymous):

\[\frac{ 4m(m+4)(m-5) }{ (m-2)(m-5) }\]

OpenStudy (anonymous):

\[\frac{ 4m(m+4) }{ (m-2) }\]

OpenStudy (anonymous):

the m+4 on top cancels out & becomes 2 \[4m(m+2)?\]

OpenStudy (anonymous):

over 1?

hartnn (hartnn):

nothing can get cancelled from \(\frac{ 4m(m+4) }{ (m-2) }\) thats your final answer

OpenStudy (anonymous):

oooh only if the 2 & the 4 were in opposite places...got it!

hartnn (hartnn):

not even then :P